OpenStudy (anonymous):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

OpenStudy (anonymous):

x=-14/11

OpenStudy (the_fizicx99):

@RobertoTB Stop giving out answers.

OpenStudy (anonymous):

We're going to use the distributive property a bunch of times here. 3(x -2) +4(2x -6) = 6(x -4) +8(2x +1) Let's break it down by each term. 3(x -2) can be rewritten as x*3 -2*3. This simplifies to 3x -6. 4(2x -6) can be rewritten as 2x*4 -6*4. This simplifies to 8x -24. 6(x -4) can be rewritten as x*6 -4*6. This simplifies to 6x -24. 8(2x +1) can be rewritten as 2x*8 +1*8. This simplifies to 16x +8. Let's put these back into the original equation. 3x -6 +8x -24 = 6x -24 +16x +8 Now, we can combine like terms (the ones with the same variable attached; in this case, either x or no variable). -3x -6 +8x -24 = 6x -24 +16x +8 -3x +8x -6 -24 = 6x +16x -24 +8 5x -30 = 22x -16 Now, you can use regular algebra to isolate x. ("Do the opposite to the opposite!")